OpenStudy (anonymous):

Convert -245° into radians

OpenStudy (anonymous):

multiply it by pi/180

OpenStudy (anonymous):

i got -4.25?

OpenStudy (anonymous):

Keep it as a frac :)

OpenStudy (anonymous):

how @FutureMathProfessor

OpenStudy (anonymous):

Just leave it as a frac and have pi left in it

OpenStudy (anonymous):

how do i convert it to a frac @FutureMathProfessor

OpenStudy (dumbcow):

\[\frac{245}{180} = \frac{49*5}{36*5} = \frac{49}{36}\]

OpenStudy (anonymous):

thank you @dumbcow

OpenStudy (anonymous):

wouldnt it be negative tho ? @dumbcow

OpenStudy (dumbcow):

yes and pi would still be there ... i was only showing how to reduce the fraction

OpenStudy (anonymous):

Given tan B = 15/9 find angle B in degrees. Round your answer to the nearest hundredth.

OpenStudy (dumbcow):

use inverse tangent button on calculator \[B = \tan^{-1} \frac{15}{9}\] calculator will give answer for 1st quadrant, to get 3rd quadrant add 180 degrees